• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 22:01:39 on January29,2008

    Entry number 227953

    keyword=Cavity Signals from 3582

    The attached figures show the cavity signals from run 3582 taken at
    20 uA during owl shift today. The pre-gain was set high for low
    currents, so somewhat above 20 uA the signals begin to saturate.

    In figure 1, the cavity position ADC value is plotted vs BPM position.


    FIGURE 1

    FIGURE 2